# Capturing Stable HDR Videos Using a Dual-Camera System

### v1

@@ III-A Dual-Camera System
Video Capture. Both cameras are triggered to start capturing images simultaneously, but subsequent image capture can occur asynchronously. One camera captures the medium-exposure reference sequence $L_{reference}(t)$, while the other captures the alternating exposure non-reference sequence $L_{non-reference}(t)$ with low and high exposure. Both cameras are set to the same resolution of $w\times h$. Due to different exposure time settings, the videos captured by the two cameras are not timestamp-synchronized. However, our method does not require precise synchronization. We use timestamp information to roughly group the data, where each group consists of frames from $L_{reference}(t)$ and two high- and low-exposure images from $L_{non-reference}(t)$ (e.g., with exposure times of (-2, 0, 2 EV), every four frames from the reference sequence, along with one pair of high- and low-exposure images, form a single group). Each frame, along with the two high- and low-exposure images, is used as input to the network. The final reconstructed HDR video matches the frame rate of $L_{reference}(t)$. To ensure a fair comparison, we capture test datasets for both AE and MEF using our system. Detailed information can be found in Sec. V-B, and the datasets will be available.
@@
Essentially, our system can be viewed as a multi-exposure HDR fusion task for dynamic scenes, where luminance alignment between neighboring frames is unnecessary due to consistently exposed reference frames. Existing MEF methods can be seamlessly extended to achieve consistent exposure results. To address challenges from video data and the variability of real-world scenes, we propose SAFNet, a more robust network, detailed in Section IV.
@@
Figure 3: The architecture of EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a restoration subnetwork. We introduce GLA and EFSM to leverage exposure information, explore the intrinsic properties of the images, and help preserve finer details across varying exposures. The asymmetric cross-feature fusion subnetwork improves image fusion by aligning cross-scale features and performing cross-feature fusion. The reconstruction subnetwork adopts a DWT-based multiscale architecture to reduce ghosting and refine features at different resolutions.
III-B Discussion.
The proposed DCS offers several advantages over traditional single-camera AE methods for HDR video reconstruction:
Stable Lighting Conditions. The DCS provides more stable exposure across frames, avoiding the fluctuations typical in alternating exposure systems.
Potential in Low-Light Conditions. The DCS performs better in low-light environments, where long exposure times in a single-camera setup can lead to artifacts and delays. The dual-camera approach mitigates these issues by distributing exposure tasks between the cameras.
Offline HDR Video Reconstruction. The DCS setup supports offline HDR video reconstruction, allowing continuous high-quality video capture without real-time processing, ensuring high-quality capture without real-time processing delays. In contrast, AE methods require real-time reconstruction to avoid flickering, which can introduce undesirable artifacts if not processed immediately.
Independent Automatic Exposure. Each camera in the DCS can independently apply established automatic exposure strategies, eliminating the need for recalibration with each frame and simplifying real-time operation. In contrast, AE methods require constant adjustments to exposure settings, complicating real-time capture.
Seamless Integration with Encoding Strategies. The dual-camera framework integrates easily with existing video encoding and decoding strategies. The secondary camera can capture frames at a low frame rate (e.g., 1fps or lower), minimizing additional bandwidth requirements, thus ensuring efficient video transmission.
@@ IV-A Overview
Given a set of LDR images of a dynamic scene $\left\{L_{1},L_{2},\ldots,L_{n}\right\}$ with different exposures, HDR imaging aims to generate an HDR image $H$ aligned to a prescribed reference image. Fig. 3 shows an overview of the proposed EAFNet. Our proposed EAFNet consists of three main subnetworks: the pre-alignment subnetwork, the asymmetric cross-feature fusion subnetwork, and the restoration subnetwork.
IV-B Pre-alignment
The pre-alignment subnetwork aims to extract meaningful details from the non-reference image by leveraging exposure information. Global luminance alignment (GLA) is introduced to maintain consistent luminance across varying exposures, while the exposure-guide feature selection module (EFSM) explores the correlation between inputs and exposure to focus on the most relevant regions for fusion.
Global Luminance Alignment. Following [53, 36], we employ three LDR images $\left\{L_{l},L_{m},L_{h}\right\}$ (sorted by their exposure length) as input. The images first undergo gamma correction [53], defined as:
$H_{i}=L_{i}^{\gamma}/{t_{i}},\ \ i=l,m,h$ (1)
where $\gamma>1$ denotes the gamma correction parameter and $t_{i}$ denotes the exposure time. The indices $l,m,h$ correspond to low, mid, and high exposure settings, respectively, with $m$ serving as the reference image. In this work, the default gamma parameter $\gamma$ is set to 2.2. Then, we will get a corresponding set of $\left\{H_{i}\right\}$.
Due to varying luminance levels at different exposure times, captured images may exhibit color rendering deviations from real-world colors, with low-exposure images losing shadow details and high-exposure images losing highlight details. Therefore, variations in exposure times certainly affect the overall color accuracy of the input LDR images. To address this, we propose a GLA in the pre-alignment subnetwork. GLA aims to correct the luminance of LDR images in the sRGB domain, producing consistent brightness features across varying exposures and enhancing detail preservation. Specifically, the LDR images are mapped to the luminance range of the reference image and achieve global luminance alignment:
$G_{i}=clip(L_{i}*(mean(L_{i})/mean(L_{m})),\ i=l,h)$ (2)
where $G_{i}$ denotes the corresponding luminance-aligned output, $mean(\cdot)$ denotes the mean luminance value, and $clip$ is a function that keeps the output in the range [0, 1]. Notably, since $L_{m}$ is the reference image, $G_{m}=L_{m}$.
@@ IV-B Pre-alignment
As shown in Fig. 4, given a set of LDR images with varying exposure times, the non-reference images can be aligned to the reference image using GLA. Finally, we can obtain 3-pairs of 9-channel tensor $L_{in}=\left(\{L_{l},H_{l},G_{l}\},\{L_{m},H_{m},G_{m}\},\{L_{h},H_{h},G_{h}\}\right)$ as the input to feed into the EFSM module.
Exposure-Guided Feature Selection Module. Existing methods primarily focus on texture for fusion, overlooking the crucial role of exposure time information, which may result in suboptimal fusion and color inconsistencies. To overcome this limitation, we propose an EFSM to explore exposure information, enhancing meaningful details and ensuring that the most relevant features are preserved despite varying exposure levels.
For a set of LDR images with corresponding exposure times of $\left\{{{t_{l}},{t_{m}},{t_{h}}}\right\}$, we standardize and relativize the input exposure times using the following formulation:
${e_{{}_{l}}}={\log_{2}}\left({{{t_{l}}}/{{t_{m}}}}\right)\cdot c,\ \ \ {e_{{}_{h}}}={\log_{2}}\left({{{t_{h}}}/{{t_{m}}}}\right)\cdot c$ (3)
@@
where ${e_{{}_{l}}}$ and ${e_{{}_{h}}}$ denote the relative exposure inputs of the low exposure branch and high exposure branch, respectively. $c$ is the hyper-parameter set to 0.1.
As shown in Fig. 4, we adopt a multiscale architecture to progressively refine feature representations across different exposure levels. The framework initiates with multi-level feature extraction: given input $L_{in}$, three parallel $3\times 3$ convolutional layers generate initial feature maps $\{f_{l},f_{m},f_{h}\}$ corresponding to low-, mid-, and high-exposure features, respectively. For subsequent processing clarity, we formally define $f^{nr}\triangleq\{f_{l},f_{h}\}$ as non-reference features and $f^{r}\triangleq f_{m}$ as reference features. Assuming there are $N$ scales in total, the reference and non-reference features of the $i$-th scales are denoted as $f_{i}^{r}$ and $f_{i}^{nr}$, respectively ($i=1$ denotes the original scale), $f_{i}^{nr}$ are fed to the shared weights exposure-guided feature selection block (EFSB) and $f_{i}^{r}$ are fed to the shared weights feature selection block (FSB). Specifically, the features undergo global average pooling, followed by two fully connected layers and a sigmoid activation to produce the feature-guide modulation coefficients $v^{nr}\in\mathbb{R}^{1\times 1\times 64}$ and $v^{r}\in\mathbb{R}^{1\times 1\times 64}$. Simultaneously, relative exposure inputs $e$ are each passed through fully connected layers, resulting in the exposure-guide modulation coefficients $v^{t}\in\mathbb{R}^{1\times 1\times 128}$.
To dynamically select and refine features from multiple input channels, the exposure-guide modulation coefficients $u^{nr}\in\mathbb{R}^{1\times 1\times 64}$ are computed as follows:
$u^{nr}=\sigma\left(FC_{3}\left(<v^{nr},v^{t}>\right)\right)$ (4)
where <,><,>< , > denotes concatenation, $\sigma$ represents sigmoid activation, $FC_{3}$ denotes three fully connected layers. The final output of EFSM can be obtained as:
$\hat{f_{i}^{r}}=f_{i}^{r}\cdot v^{r},\ \ \ \ \ \ \hat{f_{i}^{nr}}=f_{i}^{nr}\cdot u^{nr}$ (5)
IV-C Asymmetric Cross-feature Fusion
While the non-reference image may provide more details, it also poses a higher risk of introducing ghosting artifacts, especially in open scenes. To address this, we propose an asymmetric cross-attention block, which enhances the contribution of the reference input during attention computation and mitigates the potential for ghosting artifacts.
Figure 5: The structure of asymmetric cross-feature block. It employs asymmetric cross-attention to align reference and non-reference features, integrating coarse-to-fine guidance for enhanced reference-dominated feature fusion and improved HDR reconstruction.
Generally, we propose a multi-scale strategy to calculate the attention in a coarse-to-fine manner. At each scale, asymmetric cross-attention (ACA) is utilized to compute local attention. Specifically, as shown in Fig. 5, the proposed ACA first utilizes the non-reference feature as an agent, facilitating the exchange of information between the patch tokens of the reference feature and then back project it to its non-reference branch. To prioritize mid-exposure image features and encourage more effective exploration of feature interactions across different exposures, we introduce the self-information during the exchange process. Following [54], the attention matrix is thus computed as follows:
$Q_{i}=\phi(\hat{f}_{i}^{r})W_{i}^{q},\ K_{i}=\phi(\hat{f}_{i}^{nr})W_{i}^{k},\ V_{i}=\phi(\hat{f}_{i}^{nr})W_{i}^{v}$ (6)
$A_{i}=softmax((Q_{i}+K_{i})K_{i}^{T}/\sqrt{C})$ (7)
where $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{C\times C}$ are learnable weight matrices of the $i$-th scales, $\phi(\cdot)\in\mathbb{R}^{N\times C}$ denotes the unfolding operation ($N$ refers to the number of patches). Then, the aligned non-reference features $f^{anr}_{i}$ are obtained by applying a folding operation to the result of the multiplication of $A_{i}$ and $V_{i}$.
To leverage attention from coarser scales and local details at the current scale, we introduce the cross-scale guidance feature $f^{guid}$ to achieve coarse-to-fine feature fusion and interaction:
@@ IV-C Asymmetric Cross-feature Fusion
where $f^{guid}_{i}$ represent guidance features at $i$-th scales, $U$ represents pixel shuffle upsampling, $C_{v}$ denotes a $3\times 3$ convolution. As shown in Fig. 5, $f^{guid}_{i}$ learn reference-dominated attention from the previous scale, facilitating the alignment of features across scales. The output of asymmetric cross-attention block can be expressed as:
$\tilde{f}^{nr}=C_{v}(<f^{anr}_{1},U(f^{anr}_{2}),f^{guid}_{1}>)$ (9)
@@
Finally, we generate a spatial attention map with $\tilde{f}^{nr}$ and $f^{r}$ and get the final output:
$f^{out}=<\sigma(C_{v}(Relu\left(C_{v}\left(<f^{r},\tilde{f}^{nr}>\right)\right)))\cdot\tilde{f}^{nr},f_{m}>$ (10)
IV-D Restoration
The restoration subnetwork adopts a frequency-based approach to effectively mitigate ghosting artifacts in HDR images. As shown in Fig. 3, the subnetwork is designed with a multiscale architecture that progressively refines features at different resolutions. The architecture particularly emphasizes feature correction at lower scales, as these scales capture finer high-frequency details, such as textures and small features, which are essential for addressing the residual ghosting artifacts.
At the lower scales, the features undergo Discrete Wavelet Transform (DWT) to decompose the image into multi-resolution subbands. Following this, multiple Learnable Bandpass Filter (LBF) modules [55, 39] are applied to correct each frequency component. The output is then passed through the Inverse Wavelet Transform (IWT) to reconstruct the feature maps $f_{\times 1}$ in the feature domain. At higher scales, after applying DWT, a $1\times 1$ convolution is used instead of multiple LBF modules. The feature maps at higher scales primarily contain global structural information, reducing the need for complex high-frequency corrections.
Finally, we combine the high-frequency details from lower scales with the global information from higher scales, leading to improved HDR image restoration with reduced ghosting artifacts:
$H_{out}=\sigma(Conv(U(U(f_{\times 4})+f_{\times 2})+f_{\times 1}))$ (11)
where $U$ denotes upsampling, $f_{\times 4},f_{\times 2},f_{\times 1}$ are the output features at different scales, $H_{out}$ represents the final HDR image output.

### v2

@@ III-A Dual-Camera System
Video Capture and System Design. Our dual-camera system employs two identical cameras with resolution $w\times h$. In practical operation, the cameras are run asynchronously: one continuously records a medium-exposure reference sequence $L_{ref}(t_{1})$ to provide temporally consistent baseline frames, while the other alternates between low and high exposures $L_{non\mbox{-}ref}(t_{2})$ to supply the luminance diversity required for dynamic range expansion. To leverage the information from two sequences, we construct approximate pairings using timestamp metadata: as shown in Fig. 2(a), for each low/high exposure pair in the alternating stream, we gather reference frames falling within its defined temporal neighborhood and combine each of them with the same low/high pair to form input groups, enabling the reuse of one low/high pair across multiple reference frames. This pairing does not require frame-level precision, as the scenes are inherently dynamic and minor temporal offsets have a negligible impact on multi-exposure fusion. Each input group is then fed into the network to reconstruct HDR video at the same frame rate as $L_{ref}(t_{1})$.
@@
Overall, the design formulates the task as multi-exposure HDR fusion for dynamic scenes, leveraging the consistent exposure of the reference sequence to eliminate the need for luminance alignment between neighboring frames. Existing HDR image deghosting methods can be seamlessly extended to achieve consistent luminance results. To address challenges from video data and the variability of real-world scenes, we propose EAFNet, a more robust network, detailed in Section IV.
III-B Further Discussion
To further clarify our contributions, we discuss the proposed paradigm from three perspectives:
III-B1 Relationship with Previous HDR video pipelines
Traditional HDR video generation methods often rely on specialized sensors or expensive camera setups, such as Quad-Bayer sensors [20], Beam Splitter systems [28], etc. These methods produce high-quality HDR output through specialized hardware designs for capturing frames with different exposures, but come with significant hardware costs and are typically limited to high-end applications. In contrast, our dual-stream paradigm not only lowers hardware costs but also enhances flexibility, making it feasible for HDR video generation on consumer-grade devices.
Additionally, many existing methods address hardware limitations by employing alternating exposure techniques [54, 2, 8], where a single consumer-grade camera alternates between different exposure levels to capture HDR video. Our paradigm retains the cost-effectiveness of traditional AE methods while also addressing key challenges such as temporal consistency and robustness in real-world dynamic imaging scenarios.
@@ III-B1 Relationship with Previous HDR video pipelines
Figure 3: The architecture of EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a restoration subnetwork. We introduce GLA and EFSM to leverage exposure information, explore the intrinsic properties of the images, and help preserve finer details across varying exposures. The asymmetric cross-feature fusion subnetwork improves image fusion by aligning cross-scale features and performing cross-feature fusion. The reconstruction subnetwork adopts a multi-scale architecture to reduce ghosting and refine features at different resolutions.
III-B2 Relationship with Previous dual-camera systems
Dual-camera systems have been used in HDR imaging [11, 31], but they typically require strict synchronization, which is challenging in dynamic video scenes due to exposure differences between cameras, leading to ghosting artifacts. In contrast, our system performs exposure fusion without synchronization, enabling better adaptability to dynamic environments. Additionally, while synchronized systems are constrained by long exposure times that lower frame rates, our system decouples temporal stability from exposure variation, allowing flexible exposure settings and maintaining high frame rates. The reference stream controls the frame rate, making our system particularly suited for high-frame-rate HDR video capture, especially in low-light conditions.
Beyond HDR imaging, dual-camera frameworks have also been explored in related video enhancement tasks. For instance, Cheng et al. [5] and H2-Stereo [6] proposed dual-camera super-resolution systems that pair a high-resolution low-frame-rate stream with a low-resolution high-frame-rate stream to reconstruct videos with high spatiotemporal resolution. Similarly, dual-camera designs have been applied in video de-moiré [49]. While these methods also exploit asymmetric information across cameras, their optimization targets are fundamentally different from ours: they generally assume consistent exposure between cameras and focus on improving spatial or temporal fidelity, rather than expanding dynamic range. In contrast, our framework introduces an exposure-asymmetric dual-camera configuration: a reference camera maintains a stable mid-level exposure to anchor temporal consistency, while an auxiliary camera alternates between low and high exposures to provide extreme luminance details. This design effectively decouples temporal stability from exposure diversity, a direction not explored in prior work. Furthermore, we relax the requirement for frame-level synchronization by adopting group-wise asynchronous fusion, where multiple reference frames are paired with an alternating-exposure pair from the auxiliary camera and jointly processed by a dedicated deghosting network. This pipeline enables high-frame-rate HDR video capture with minimal flicker, a capability not achievable with single-camera alternating exposure or conventional dual-camera systems designed for other tasks.
In addition, our system exhibits high scalability and strong generalization capability. It does not rely on device-specific calibration or training data. Instead, multi-exposure fusion models trained on publicly available datasets can be seamlessly integrated into our framework without any retraining or fine-tuning. Leveraging this flexibility, our system can directly produce flicker-free HDR videos from real-world captures. In contrast, existing dual-camera HDR solutions typically require carefully calibrated camera pairs and task-specific training, which significantly restricts their deployment in practical scenarios. To the best of our knowledge, none of the prior methods provide such plug-and-play adaptability while maintaining high temporal stability.
III-B3 The advantages of the proposed video generation paradigm
The proposed video generation paradigm offers several distinct advantages. First, the dual-stream design introduces an exposure-asymmetric configuration that explicitly decouples temporal stability from exposure diversity. This principle alleviates the feature interference inherent in AE-based HDR systems and improves robustness in dynamic real-world environments.
Second, the paradigm offers flexibility in exposure scheduling and frame rate control. By anchoring the output frame rate to the reference stream, the system can deliver high-frame-rate HDR video while simultaneously leveraging the auxiliary stream for extended dynamic range, which is particularly beneficial in low-light or high-motion conditions. In this asynchronous dual-stream capture scheme, the reference stream follows standard auto-exposure to provide a stable baseline, while the auxiliary stream operates at a lower frame rate with fixed or preset exposures. Such a design improves energy efficiency and reduces bandwidth and storage overhead, enabling seamless integration with standard video encoding and transmission protocols.
Third, the paradigm achieves strong scalability and generalization. Multi-exposure fusion models trained on publicly available datasets can be directly applied without device-specific calibration or retraining, enabling plug-and-play deployment.
Finally, the design is highly practical for integration into consumer-grade devices. Existing multi-camera systems can readily adopt the dual-stream setup without the need for specialized sensors or costly calibration procedures, broadening the accessibility of HDR video capture.
@@ IV-A Overview
In this section, we introduce the overall framework of the proposed EAFNet, built upon a pre-alignment subnetwork (in Sec. IV-B), an asymmetric cross-feature fusion subnetwork (in Sec. IV-C) and a restoration subnetwork (in Sec. IV-D).
As shown in Fig. 3, the pre-alignment subnetwork aims to improve detail preservation and color accuracy by exploring exposure information and extracting meaningful features from the non-reference image. It incorporates Global Luminance Alignment (GLA) to ensure consistent brightness across varying exposures, and an Exposure-guide Feature Selection Module (EFSM) to identify and focus on the most relevant regions based on the exposure information. These mechanisms work together to enhance the input features before fusion. Next, the asymmetric cross-feature fusion subnetwork plays a critical role in blending features from different exposures while emphasizing the most important details. It processes the low- and high-exposure inputs through separate asymmetric cross-feature fusion blocks, each built with a multi-scale design. The fusion is driven by an Asymmetric Cross-Attention (ACA), which leverages reference-dominated attention maps to improve the alignment of query and key features, facilitating more effective feature propagation across multiple scales. Finally, the restoration subnetwork addresses common issues such as ghosting artifacts. It is designed to suppress ghosting while preserving fine details at multiple scales.
IV-B pre-alignment subnetwork
The pre-alignment subnetwork aims to extract meaningful details from the non-reference image by leveraging exposure information. GLA is introduced to maintain consistent luminance across varying exposures, while the EFSM explores the correlation between inputs and exposure to focus on the most relevant regions for fusion.
Global Luminance Alignment (GLA). Following established HDR reconstruction pipelines [23, 56], we employ three LDR images $\{{L}_{l},{L}_{m},{L}_{h}\}$ (sorted by their exposure length) as input. These images are first mapped to the HDR domain relying on gamma correction [23] to produce a set of gamma-corrected images ${G}_{i}$}, where $i\in\{l,m,h\}$ and $\gamma=2.2$.
To reduce inter-exposure luminance discrepancies and improve alignment robustness, we introduce GLA in the pre-alignment stage. GLA operates in the sRGB domain and ensures consistent luminance distributions across inputs and minimizes exposure-induced mismatches during fusion. Formally, for each non-reference LDR image $L_{i}$, global alignment is performed as:
$A_{i}=\operatorname{clip}\!\left(L_{i}\times\frac{\mathbb{E}(L_{m})}{\mathbb{E}(L_{i})},\;0,1\right),\quad i\in\{l,h\},$ (1)
where $A_{i}$ denotes the luminance-aligned image, $\mathbb{E}(\cdot)$ computes the average luminance, and $\operatorname{clip}(\cdot,0,1)$ restricts pixel values to the valid dynamic range. Since the mid-exposure image $L_{m}$ serves as the reference, we set $A_{m}=L_{m}$.
@@
As shown in Fig. 3, given a set of LDR images with varying exposure times, we construct the input tensor $\mathbf{L}_{\text{in}}=\big{(}\{L_{l},G_{l},A_{l}\},\{L_{m},G_{m},A_{m}\},\{L_{h},G_{h},A_{h}\}\big{)}$, where each triplet $(L_{i},G_{i},A_{i})$ represents the original, gamma-corrected, and GLA-processed versions of the image under exposure setting $i\in\{l,m,h\}$. This tensor is then passed to the EFSM for feature selection and alignment refinement.
Exposure-guided Feature Selection Module (EFSM). Exposure information is crucial for distinguishing the reliability and detail across different exposures. For the obtained multi-channel features, we propose the EFSM, which consists of the Exposure-guided Feature Selection Block (EFSB) and the Feature Selection Block (FSB). EFSM introduces exposure-aware modulation based on relative exposure priors, selectively emphasizing well-exposed regions and suppressing unreliable features, thereby ensuring more reliable feature representation for fusion.
Specifically, EFSM operates in a multi-scale feature refinement architecture, as depicted in Fig. 4. Given a 9-channel composite input tensor $\mathbf{L}_{\text{in}}$, we first extract three initial feature maps $\{\mathbf{F}_{l},\mathbf{F}_{m},\mathbf{F}_{h}\}$ via parallel $3\times 3$ convolutional layers. For subsequent processing clarity, we formally define $\mathbf{F}^{nr}\triangleq\{\mathbf{F}_{l},\mathbf{F}_{h}\}$ as non-reference features and $\mathbf{F}^{r}\triangleq\mathbf{F}_{m}$ as reference features. These are fed into two distinct but weight-shared selection blocks: EFSB for non-reference features and FSB for reference features.
To encode exposure priors, we normalize the exposure times $\{t_{l},t_{m},t_{h}\}$ relative to the middle exposure as:
$e_{l}=\log_{2}\!\left(\frac{t_{l}}{t_{m}}\right)\cdot c,\quad e_{h}=\log_{2}\!\left(\frac{t_{h}}{t_{m}}\right)\cdot c,$ (2)
@@
where $e_{l}$ and $e_{h}$ represent the relative exposure embeddings for the low- and high-exposure branches, respectively, and $c$ is a hyper-parameter empirically set to $0.1$.
@@
$\mathbf{u}^{\text{nr}}=\sigma\!\big{(}\operatorname{FC}_{3}([\mathbf{v}^{\text{nr}};\mathbf{v}^{\text{t}}])\big{)},$ (3)
where $\cdot$$\cdot$] denotes channel-wise concatenation, $\sigma(\cdot)$ is sigmoid activation, and $\operatorname{FC}_{3}(\cdot)$ represents three fully connected layers. Finally, the refined features are obtained by reweighting:
$\hat{\mathbf{F}}_{i}^{\text{r}}=\mathbf{F}_{i}^{\text{r}}\odot\mathbf{v}^{\text{r}},\quad\hat{\mathbf{F}}_{i}^{\text{nr}}=\mathbf{F}_{i}^{\text{nr}}\odot\mathbf{u}^{\text{nr}},$ (4)
where $\odot$ denotes element-wise multiplication.
IV-C asymmetric cross-feature fusion subnetwork
One of the central challenges in multi-exposure fusion lies in leveraging the complementary details from non-reference images without introducing ghosting artifacts, particularly in the presence of large motion or occlusions [58, 34]. While non-reference inputs offer richer information in under-/over-exposed regions, they also carry a higher risk of misalignment. To address this, we propose an asymmetric cross-feature fusion subnetwork, which employs asymmetric cross-feature fusion blocks to selectively align and fuse cross-exposure features in a coarse-to-fine manner. Each fusion block is centered around the Asymmetric Cross-Attention (ACA) module, which leverages reference-dominant attention to fuse features.
Figure 5: The structure of the asymmetric cross-feature fusion block. It employs asymmetric cross-feature attention to align reference and non-reference features, integrating coarse-to-fine guidance for enhanced reference-dominated feature fusion and improved HDR reconstruction.
As illustrated in Fig. 5, the proposed module adopts a multi-scale architecture, where local attention at each level is computed via an ACA mechanism. At each scale $i\in\{1,2,3\}$, we first unfold the refined reference features $\hat{\mathbf{F}}_{i}^{\text{r}}$ and non-reference features $\hat{\mathbf{F}}_{i}^{\text{nr}}$ into patch tokens using a window-based unfolding operator $\phi(\cdot)$. The attention is formulated in an asymmetric manner, where queries are derived from the reference stream, while keys and values are obtained from the non-reference stream to facilitate reference-guided feature aggregation. To further emphasize reference dominance, we incorporate query features into the key projection via additive modulation:
$\mathbf{M}_{i}=\operatorname{softmax}\!\left((\mathbf{Q}_{i}+\mathbf{K}_{i})\mathbf{K}_{i}^{\top}/{\sqrt{C}}\right),$ (5)
where $\mathbf{Q}_{i}=\phi(\hat{\mathbf{F}}_{i}^{\text{r}})\,\mathbf{W}_{i}^{q}$ and $\mathbf{K}_{i},\mathbf{V}_{i}=\phi(\hat{\mathbf{F}}_{i}^{\text{nr}})\,\mathbf{W}_{i}^{k,v}$, with $\mathbf{W}_{i}^{q,k,v}$ being learnable matrices. This asymmetric formulation encourages stronger alignment toward the reference features and reduces the impact of unreliable non-reference content. Then, the aligned non-reference features $\mathbf{F}^{anr}_{i}$ are obtained by applying a folding operation to the result of the multiplication of $\mathbf{M}_{i}$ and $\mathbf{V}_{i}$.
To enhance fine-scale alignment with global context, we incorporate a cross-scale guidance mechanism. Specifically, attention information from the coarser scale $i+1$ is projected and injected into the finer scale $i$ as:
$\mathbf{F}_{i}^{\text{guid}}=\operatorname{Conv}_{3\times 3}\big{(}[\mathbf{F}_{i}^{\text{nr}};U(\psi(\mathbf{M}_{i+1}))]\big{)},\quad 1\leq i<N,$ (6)
where $\operatorname{Conv}_{3\times 3}(\cdot)$ denotes a $3\times 3$ convolution, $N$ denotes the number of refinement scales, $U(\cdot)$ is pixel shuffle upsampling, and $\psi(\cdot)$ represents the projection of attention from the coarser scale. Then, $\mathbf{F}_{i}^{\text{guid}}$ injects reference-dominated attention to facilitate multi-scale alignment. The output of the ACA $\mathbf{F}^{\text{nr}}$ is computed by merging the aligned features from multiple scales, combining the finer-scale features with guidance from the coarser scales to produce a refined feature representation.
@@
Finally, we generate a spatial attention map using $\tilde{\mathbf{F}}^{\text{nr}}$ and $\mathbf{F}^{\text{ref}}$ and get the fusion output.
IV-D restoration subnetwork
As shown in Fig. 3, the restoration subnetwork is designed to suppress ghosting artifacts while preserving fine details across multiple scales.
Each scale begins with a Discrete Wavelet Transform (DWT) to decompose features into frequency subbands. At the lowest scale, where high-frequency details such as textures and residual motion artifacts are most prominent, we apply a Lightweight DomainPlus Block (LDPB) to perform frequency-specific corrections. The LDPB is adapted from [62] by removing dense connections to reduce complexity while retaining frequency modeling capacity.
In contrast, higher scales capture more global and low-frequency content. These layers are processed using a simple $1\times 1$ convolution after DWT, which is sufficient for coarse structure refinement without the need for full spectral filtering. All features are subsequently reconstructed using the Inverse Wavelet Transform (IWT).
The refined features from each scale are fused in a top-down manner using upsampling and residual connection to produce the final HDR output.